OpenStudy (anonymous):

Simplify Using All Positive Exponents: (x^-1u^2/3z^2)^-3(uz^5)

OpenStudy (anonymous):

\[(x^-1u^2/3z^2)^-3(uz^5)\] \[={x^3*u^-6*z^6}{27} *uz^5\] \[=\frac{27x^3z^11}{u^5}\] =(27 x^3 z^11)/u^5

OpenStudy (anonymous):

Can You help Me With Another Problem Please